OpenStudy (anonymous):

what is the no. of moles of radiation energy absorbed? If energy of absorption is 6e-7 erg & wavelength is 3600 angstrom

OpenStudy (aaronq):

The energy of 1 photon of 3600 angstrom (360 nm) wavelength is given by: \(E_{photon}=\dfrac{hc}{\lambda}=\dfrac{{(6.626*10^{-34}~m^2~kg/s})*({3*10^8~m/s})}{(360*10^{-9}m)}\) \(=5.52*10^{-19}~J\) Im guessing that "6e-7 erg" means \(6.0*10^{-7}~J\). Divide this energy by the energy of 1 photon to find how many photons were absorbed. Convert this number to moles by dividing it by Avogadro's number.